computer projections tell you that an investment that compounds monthly will yield a future value of FV = PV (1.07763) ^ t, where t is in years.

What is the rate, r, for this account? since we have two expressions for FV here, try starting with the following :

FV = PV(1 + -) 12+ = PV(1.07763)

which after dividing by PV and using the laws of exponents, can be rewritten

(1+)) = 10703)

Now use the fact that 19 = implies that a = b, to find r.
Answer #1

1. 0776.3 Com 3)* = (1.01163) (+)= (1.07763) - 1 + 1/2 = 125 1.07763 71+ F = 1.00624 = 1.00624 şi = 0.00 G 2.4 str.= 0.0 748